OpenStudy (anonymous):

Calculus? apply the mean value theorem to f on [0,1] to find all values of c such that f'(c)= f(b)-f(a)/b-a. f(x)=x^3

ganeshie8 (ganeshie8):

start by evaluating f(a), f(b) and f'(x)

ganeshie8 (ganeshie8):

f(a) = ? f(b) = ? f'(x) = ?

OpenStudy (anonymous):

f'(x)=3x^2

OpenStudy (anonymous):

a is 0 and b is 1 right?

ganeshie8 (ganeshie8):

yup! a and b are just the boundaries in given interval

ganeshie8 (ganeshie8):

f'(x)=3x^2 f'(c) = 3c^2

ganeshie8 (ganeshie8):

a = 0, b = 1 f(0) = 0^3 = 0 f(1) = 1^3 = 1

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

f'(c)= f(b)-f(a)/b-a plug them above

ganeshie8 (ganeshie8):

3c^2 = 1-0/1-0 3c^2 = 1/1 3c^2 = 1 c^2 = 1/3 c = ?

OpenStudy (anonymous):

c=0.576 ?

OpenStudy (anonymous):

cause you square both sides right?

ganeshie8 (ganeshie8):

c = 1/sqrt(3) = 0.576 is \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

notice that we discarded the negative value -1/sqrt(3) cuz, it doesnt lie in given interval : [0, 1]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

again thank you so much

ganeshie8 (ganeshie8):

hey wait a sec, we did a mistake

OpenStudy (anonymous):

where?

ganeshie8 (ganeshie8):

nope, it looks perfect ! ok

ganeshie8 (ganeshie8):

;)

OpenStudy (anonymous):

lol thank you ;)

ganeshie8 (ganeshie8):

np, u wlc :)